# Geometric Autoencoder Priors for Bayesian Inversion: Learn First Observe Later

### v1

@@ 2.3 Observational Noise Estimation
Beyond the benefits of quantifying uncertainty about model predictions, the Bayesian paradigm offers a principled framework for estimating other aspects of the data-generating model. We highlight that such estimations are completely independent of the training of the autoencoder prior – this is not the case for direct regression methods which need to have all observation variables incorporated at training time. We consider now the common task of jointly estimating the observational noise standard deviation $\sigma$, along with the field of interest $\mathbf{u}_{n}$. This is achieved through the latent joint posterior
$p_{\mathbf{z},\sigma|\mathbf{y}_{o}}(\mathbf{z},\sigma)=\frac{p_{\mathbf{y}_{o}|\mathbf{z},\sigma}(\mathbf{y}_{o})p_{\mathbf{z},\sigma_{o}}(\mathbf{z}_{o},\sigma)}{p_{\mathbf{y}_{o}}(\mathbf{y}_{o})}.$ (8)
@@
We can sample the joint posterior on the solution fields through the pushforward $p_{\mathbf{u}_{o},\sigma_{o}|\mathbf{y}_{o}}=(D^{\psi}_{o}\otimes\mathrm{Id}.)_{\,\#}\,p_{\mathbf{z},\sigma_{o}|\mathbf{y}_{o}},$ where $\mathrm{Id}.$ is the identity map.
@@ 3 Implementation
8:return return $N_{a}$ samples of $\mathbf{u}_{n}^{(1:N_{s})}$ with least residual $r_{i}$.
@@
Sampling: To realize GABI, we first train the autoencoder; then, we sample from the corresponding latent posterior as in (6) and decode the resulting latent samples to obtain samples from the posterior for a given geometry. We note two alternatives for sampling from the posterior: MCMC, and ABC sampling. We find ABC sampling to be advantageous for this task due to: (i) the massive parallelization capabilities of neural networks as opposed to the sequential nature of MCMC; (ii) the specification of a well tailored prior distribution – the prior is not unnecessarily vague leading to good overlap between the prior and posterior; (iii) the low dimensionality of the observational data – removing the need to use possibly problematic summary statistics. In Algorithm 1 we outline the implementation of the GABI-ABC variant. For MCMC sampling, we use the NUTS algorithm (Hoffman et al., 2014).
Notes on Alternative Approaches: Physics Driven Inversion. These require additional knowledge aside from access to a representative dataset and information specific to each system we wish to query such as the governing PDE, boundary conditions, forcing etc. VAEs. We refer the reader to Appendix C for an in depth discussion on the unsuitability of VAEs for the task we are solving, both in terms of the overall methodology and in terms of the autoencoder used within the proposed methodology. Supervised learning/Direct Map Inversion. One can choose to learn a supervised learning direct mapping $(\mathbf{y}_{n},\mathsf{M}_{n})\mapsto\mathbf{u}_{n}$; this approach is the main point of comparison with our method. The significant disadvantage of this approach is that the observation process (in this paper $\mathbf{y}_{n}=\mathbf{H}_{n}\mathbf{u}_{n}+\bm{\xi}_{n}$) must be known at training time and kept the same at inference time (or kept the same in a distributional sense, as explained in the comparisons). However, in practice, the observation process for different physical systems can vary enormously as we measure different quantities such as moments, local averages, summary coefficients, have a variable number of observations, or vary the type and magnitude of noise etc. In contrast to this, GABI is independent of the observation process at training time and only requires this information at test time. Thus, GABI offers a train-once-use-everywhere type foundation model, hugely advantageous for practical/industry uptake of ML models. Bayesian NNs. The problem of specifying informative priors over varying geometries persists and inherits the problems of direct maps discussed above.
@@ 4.1 Rectangles – Steady-State Heat
In this section we study the homogenous steady-state heat equation on rectangular domains with varying width, height, and boundary conditions. Thus, for a given rectangle $\mathcal{M}_{n}$, boundary condition $h_{n}$, any $u_{n}$ satisfies the steady-state heat equation. We take $\mathcal{M}_{n}\subset\mathbb{R}^{2}$ to be $\mathcal{M}_{n}=\{x\in\mathbb{R}^{2}:0<x_{1}<l_{n},0<x_{2}<w_{n}\}$. The domain $\mathcal{M}_{n}$ is defined by its length and width $(l_{n},w_{n})$ drawn from $\texttt{uniform}((0.1,0.1),(1,1))$; we specify $h_{n}$ to be zero on the bottom and left side, and drawn from $\texttt{uniform}((0.1,0),(1,1))$ on the top and right. The solution to (9) depends on $\mathcal{M}_{n},h_{n}$. We generate 1k solutions to be in the training dataset. In Figure 1 we show example geometries and solution in the dataset. Once the model in trained on this data (the loss in shown in B.3) we select a new observation vector and geometry $\mathsf{M}_{n},\mathbf{y}_{n}$ and use ABC sampling as in Algorithm 1; we show the predictive mean and standard deviation of the sampled solution fields in Figure 6. Figure 2 shows the histogram of the predictive posterior at four query locations.
@@
Comparison – Known Noise. In Table 1 we compare two sampling variants of GABI with a direct regression map as well as Gaussian Process Kriging with various kernels [^3] GABI-NUTS was tested for $10^{2}$ runs not $10^{3}$ due to prediction runtime.; M stands for Matérn. In this setup we keep number of observations at 10 randomly selected locations and the noise standard deviation at $10^{-2}$. Comparison – Unknown Noise. In Table 2 we test the noise estimation capabilities of GABI. We have the noise standard deviation be random and drawn from a shifted log-normal distribution as $\bm{\xi}_{n}=\exp(\bm{\varepsilon}-4)+10^{-3};\bm{\varepsilon_{n}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. We estimate the noise for each data $\mathbf{y}_{n}$ with the GABI approach outlined in 2.3.
@@ 4.2 Airfoils – Flow Field
In this section we test the proposed methodology on the full field reconstruction or airflow around an airfoil from a dataset of Reynolds averaged Navier-Stokes simulations. We use the setup from Duthé et al. (2025). Here the geometric autoencoder is trained on the pressure and 2D velocity field of 1k airfoil geometries with various inflow conditions. Once trained, the observation vector $\mathbf{y}_{o}$ is taken to be noisy pressure observations at a small number of nodes on the surface of the airfoil, emulating a physical scenario when such pressure sensors may be attached to a wing/blade. The Bayesian inversion task is to reconstruct the pressure and velocity field around the airfoil, a task far too ill-posed without a highly informative prior. To compile the results in Table 3 we test GABI-ABC against the Direct Map with the same architecture type – we randomize the number of observation locations to be drawn between 5 and 50 observations with probability inversely proportional to the number of observation locations. For the supervised Direct Map approach the distribution of the number of observation must be known at train time, this is not the case for GABI who’s training is independent of the observation process. Figure 3 shows a select number of results.
@@ 4.3 Car Body – Acoustic Vibration and Source Localization
Here, we study an acoustic resonance problem with a source localization task. We have a collection of car body geometries which undergo damped vibration due to a localized Gaussian bump forcing present in the first $1/5$ of the vehicle (where the engine is), the inversion task is to reconstruct the full field acoustic vibration amplitude, $u$, as well as reconstruct the full field forcing, $f$ , on the surface of the vehicle from a small number of sparse noisy measurements of the vibration amplitude, we do not observe the forcing field. The car geometries are taken from (Umetani & Bickel, 2018). We train the model on 500 car geometries, the number of observation location is randomized as in Section 4.2. The numerical results are in Table 4 and we show a reconstruction result in 4.
@@ 4.4 Terrain – Flow Field
To demonstrate the scalability of our approach, we apply it to the reconstruction of flow fields over complex terrain. Here, we use a large dataset (127GB) of more than 5k RANS simulations, produced by Achermann et al. (2024). From the original $64^{3}$ grids, we extract 8-voxel-thick, terrain-conforming layers, resulting in $64\times 64\times 8$ subdomains. Each subdomain is then converted into a graph, where voxels become nodes connected to their adjacent neighbors. We scale both the training and the inference methods with multi-GPU pipelines. Using four RTX4090 GPUs, training takes 52hr while inference takes 480s. Figure 5 shows the GABI-ABC pressure and velocity results.
@@ Appendix A Proofs
Remember
$Z(y)=\int_{\mathcal{U}}\exp\left(-\Phi(u;y)\right)\mathrm{d}\mathbb{P}_{u};\quad\quad\mathbb{P}_{u}:=g_{\,\#}\mathbb{P}_{z}.$
@@
where second equality is given by Theorem 3.6.1 in Bogachev & Ruas (2007). Applying the same equality, $\forall A\in\mathcal{F}$,
@@
This follows directly from Lemma 2.1 and Theorem 2.2 by identifying $z$ with $(\mathbf{z},\sigma)$, $\mathbb{P}_{z}$ with $(\mathbb{P}_{\mathbf{z}}\otimes\mathbb{P}_{\sigma})$, $g$ with $(D^{\psi}_{n}\otimes\mathrm{Id}.)$, changing to potential function to $\Phi(u,\sigma;y)=\frac{1}{2\sigma^{2}}\|y-Hu\|^{2}_{2}$, and finally noting $\mathrm{Id}._{\#}\mathbb{P}_{\sigma}=\mathbb{P}_{\sigma}$. ∎
@@ Appendix B Additional Numerical Results and Details
  • Geometry aware architecture and data-structure. As the main premise of this work is to share information about physical fields which depend on the geometry of the problem to which they relate, it is essential that the data-structure reflect this geometric information, and that the neural networks used be effective at processing this information.
  • Mapping to and from fixed-dimensional vectors. This framework relies on learning latent representations between solution to PDEs and the geometries on which these are defined. Hence, one needs to map to and from fixed dimensional latent spaces from arbitrary geometries.
@@ B.1 Graph Convolutional Network
In GABI: In the encoder $E^{\theta}$ we map attributed graphs to latent vectors of fixed dimensions. The decoder maps a pairing between a latent variable and a graph and populates the node attributes on the input graph. Hence the last layer in a channel wise averaging across the graph and the channel dimension is taken to be the latent dimension. The decoder $D^{\psi}$ takes as argument a graph where the node attributes are only the coordinates, hence only contains geometric information, and we expand the latent sample $\mathbf{z}$ to decode to populate all nodes with channel dimension equal the latent dimension.
@@ B.3.2 Additional Results
In Figure 6 we show the results of inference with GABI-ABC on for the full field reconstructing based on the small number of sparse noisy observations in red. In Figure 7 we show the two terms in the loss for training the GABI autoencoder as well as a histogram of the encoded data batch overlaid with the 1D standard normal distribution. We then test the performance of ABC sampling vs NUTS sampling as we vary the dimension of the latent space for the heat inversion task. The MAE, quality of UQ, and inference times are shown in Figure 8.
@@
Figure 8: Sampling schemes on the same autoencoder model with the same channel space and different latent dimensions. We compare (a) mean absolute predictive error, (b) uncertainty quantification under a Gaussian assumption, (c) sampling time, .
@@ B.4.2 Additional Results
Figure 11: Airfoil field samples drawn from the geometry conditioned joint prior over pressure, $v_{x}$, and $v_{y}$ $(\mathbf{u}_{m}\sim D^{\psi}_{m\,\#}q_{\mathbf{z}})$ for a given geometry.

### v2

@@ 2.3 Observational Noise Estimation
Beyond the benefits of quantifying uncertainty related to model predictions, the Bayesian paradigm offers a principled framework for estimating other aspects of the data-generating model. We highlight that such estimations are completely independent of the training of the autoencoder prior – this is not the case for direct regression methods which need to have all observation variables incorporated at training time. We now consider the common task of jointly estimating the observational noise standard deviation $\sigma_{o}$ (now indexed by $o$ as it may vary from one system to another), together with the field of interest $\mathbf{u}_{o}$. This is achieved through the latent joint posterior
$p^{\psi}_{\mathbf{z},\sigma_{o}|\mathbf{y}_{o}}(\mathbf{z},\sigma_{o})=\frac{p_{\mathbf{y}_{o}|\mathbf{z},\sigma_{o}}^{\psi}(\mathbf{y}_{o})p_{\mathbf{z},\sigma_{o}}(\mathbf{z},\sigma_{o})}{p_{\mathbf{y}_{o}}(\mathbf{y}_{o})}.$ (8)
where $p_{\mathbf{z},\sigma_{o}}(\mathbf{z},\sigma_{o})=q_{\mathbf{z}}(\mathbf{z})p_{\sigma_{o}}(\sigma_{o})$ for a specified prior $p_{\sigma_{o}}(\sigma_{o})$.
@@
We can sample the joint posterior on the solution fields through the pushforward $p_{\mathbf{u}_{o},\sigma_{o}|\mathbf{y}_{o}}^{\psi}=(D^{\psi}_{o}\otimes\mathrm{Id}.)_{\,\#}\,p^{\psi}_{\mathbf{z},\sigma_{o}|\mathbf{y}_{o}},$ where $\mathrm{Id}.$ is the identity map.
@@ 3 Implementation
8:return return $N_{a}$ samples of $\mathbf{u}_{o}^{(1:N_{s})}$ with least residual $r_{i}$.
@@
Sampling: To realize GABI, we first train the autoencoder; then, we sample from the corresponding latent posterior as in (6) and decode the resulting latent samples to obtain samples from the posterior for a given geometry. We note two alternatives for sampling from the posterior: MCMC, and ABC sampling. We find ABC sampling to be advantageous for this task due to: (i) the massive parallelization capabilities of neural networks as opposed to the sequential nature of MCMC; (ii) the specification of a well-tailored prior distribution – the prior is not unnecessarily vague leading to good overlap between the prior and posterior; (iii) the low dimensionality of the observational data – removing the need to use possibly problematic summary statistics. In Algorithm 1 we outline the implementation of the GABI-ABC variant. For MCMC sampling, we use the NUTS algorithm (Hoffman et al., 2014).
Notes on Alternative Approaches: Physics Driven Inversion. These require additional knowledge aside from access to a representative dataset and information specific to each system we wish to query such as the governing PDE, boundary conditions, forcing etc. VAEs. We refer the reader to Appendix C for an in-depth discussion on the unsuitability of VAEs for the task we are solving, both in terms of the overall methodology and in terms of the autoencoder used within the proposed methodology. Supervised learning/Direct Map Inversion. One can choose to learn a supervised learning direct mapping $(\mathbf{y}_{n},\mathsf{M}_{n})\mapsto\mathbf{u}_{n}$; this approach is the main point of comparison with our method. The significant disadvantage of this approach is that the observation process (in this paper $\mathbf{y}_{n}=\mathbf{H}_{n}\mathbf{u}_{n}+\bm{\xi}_{n}$) must be known at training time and kept the same at inference time (or kept the same in a distributional sense, as explained in the comparisons). However, in practice, the observation process for different physical systems can vary enormously as we measure different quantities such as moments, local averages, summary coefficients, have a variable number of observations, or vary the type and magnitude of noise etc. In contrast to this, GABI is independent of the observation process at training time and only requires this information at test time. Thus, GABI offers a train-once-use-everywhere type foundation model, hugely advantageous for practical/industry uptake of ML models. Gaussian Process Regression. For an inverse problem over fields, a natural choice of prior is a Gaussian process prior. As the observation process $\mathbf{H}_{o}$ is linear, the classical Bayesian posterior for this problem is precisely a Gaussian process regression (GPs). We perform hyper-parameter optimization of this GP, which relates it to common usage of GPs in the context of ML. The GPs are defined directly on the graphs (Borovitskiy et al., 2021). Bayesian NNs. The problem of specifying informative priors over varying geometries persists and inherits the problems of direct maps discussed above.
@@ 4 Numerics
In this section we present numerical results for steady-state heat on rectangles, flow fields around airfoils, damped resonance and source localization on car bodies, and terrain flows. We compare our method to direct regression (Direct Map) using graph NNs, and Gaussian process regression with Matérn kernels $1/2,3/2,$ and radial basis function (RBF). In all experiments we report the wall-clock training time (Train) on Nvidia RTX 4090 GPUs (the terrain problem is multi-GPU), and the mean per-geometry prediction time (Pred.). We also note that in using ABC for sampling from the posterior we do not need to compute the likelihood function. This may have strong advantages in application where the observation process yields intractable likelihoods. In using GABI with MCMC, as in the comparisons, we do compute the likelihood. In training the latent prior for GABI, the main way we have of knowing if the autoencoder networks perform adequately for the inference task, is by assessing the quality of reconstruction for a holdout set, as well as how close the embed holdout set is to being Gaussian in the latent space, which assessed qualitatively and quantitatively.
@@ 4.1 Rectangles – Steady-State Heat
In this section, we study the homogeneous steady-state heat equation on rectangular domains with varying width, height, and boundary conditions. Thus, for a given rectangle $\mathcal{M}_{n}$, boundary condition $h_{n}$, any $u_{n}$ satisfies the steady-state heat equation. We take $\mathcal{M}_{n}\subset\mathbb{R}^{2}$ to be $\mathcal{M}_{n}=\{x\in\mathbb{R}^{2}:0<x_{1}<l_{n},0<x_{2}<w_{n}\}$. The domain $\mathcal{M}_{n}$ is defined by its length and width $(l_{n},w_{n})$ drawn from $\texttt{uniform}((0.1,0.1),(1,1))$; we specify $h_{n}$ to be zero on the bottom and left side, and drawn from $\texttt{uniform}((0.1,0),(1,1))$ on the top and right. The solution to (9) depends on $\mathcal{M}_{n},h_{n}$. We generate 1k solutions to be in the training dataset. In Figure 1 we show example geometries and solutions in the dataset. Once the model is trained on this data (the loss is shown in B.3) we select a new observation vector and geometry $\mathsf{M}_{o},\mathbf{y}_{o}$ and use ABC sampling as in Algorithm 1. Figure 2 (b) shows 4 coloured histograms corresponding to the marginal distribution of a decoded posterior (based on 10 sparse observations in the vector $\mathbf{y}_{o}$) at the 4 queried locations on Figure 2 (a). As the orange dot is close to a boundary with value 0 in the entire dataset (4 representative samples from the 1k geometry dataset can be seen in Figure 1), GABI presents very small uncertainty in this area. As we query points in areas where there is more variability of the solution in the dataset, GABI presents wider uncertainty. In all cases the mean of the GABI posterior is very close to the ground truth (vertical dashed line). This exemplifies the balance GABI strikes between sparse data and informative data-driven priors.
@@
Comparison – Known Noise. In Table 1 we compare two sampling variants of GABI with a direct regression map as well as Gaussian Process Kriging with various kernels [^3] GABI-NUTS was tested for $10^{2}$ runs not $10^{3}$ due to prediction runtime.; M stands for Matérn. In this setup, we keep the number of observations at 10 randomly selected locations and the noise standard deviation at $10^{-2}$. Comparison – Unknown Noise. In Table 2 we test the noise estimation capabilities of GABI. We have the noise standard deviation be random and drawn from a shifted log-normal distribution as $\sigma_{n}=\exp({\varepsilon_{n}}-4)+10^{-3};{\varepsilon_{n}}\sim\mathcal{N}(0,I)$. We estimate the noise for each data $\mathbf{y}_{o}$ with the GABI approach outlined in 2.3.
@@ 4.2 Airfoils – Flow Field
In this section, we test the proposed methodology on the full field reconstruction or airflow around an airfoil from a dataset of Reynolds averaged Navier-Stokes simulations. We use the setup described in Duthé et al. (2025). Here, the geometric autoencoder is trained on the pressure and 2D velocity field of 1k airfoil geometries with various inflow conditions. Once trained, the observation vector $\mathbf{y}_{o}$ is taken to be noisy pressure observations at a small number of nodes on the surface of the airfoil, emulating a physical scenario when such pressure sensors may be attached to a wing/blade. The Bayesian inversion task is to reconstruct the pressure and velocity field around the airfoil, a task far too ill-posed without a highly informative prior. To compile the results in Table 3 we test GABI-ABC against the Direct Map with the same architecture type – we randomize the number of observation locations to be drawn between 5 and 50 observations with probability inversely proportional to the number of observation locations. For the supervised Direct Map approach the distribution of the number of observation must be known at train time, this is not the case for GABI who’s training is independent of the observation process. Figure 3 shows a select number of results. For a discussion on out-of-distribution inference we refer the reader to Appendix B.4.2.
@@ 4.3 Car Body – Acoustic Vibration and Source Localization
Here, we study an acoustic resonance problem with a source localization task. We have a collection of car body geometries which undergo damped vibration due to a localized Gaussian bump forcing present in the first $1/5$ of the vehicle (where the engine is), the inversion task is to reconstruct the full field acoustic vibration amplitude, $u$, as well as the full field forcing, $f$, on the surface of the vehicle from a small number of sparse noisy measurements of the vibration amplitude. We do not observe the forcing field. The car geometries are taken from (Umetani & Bickel, 2018). We train the model on 500 car geometries, the number of observation locations is randomized as in Section 4.2. The numerical results are in Table 4 and we show a reconstruction result in Figure 4.
@@ 4.4 Terrain – Flow Field
To demonstrate the scalability of our approach, we apply it to the reconstruction of flow fields over complex terrain. Here, we use a large dataset (127GB) of more than 5k RANS simulations, produced by Achermann et al. (2024). From the original $64^{3}$ grids, we extract 8-voxel-thick, terrain-conforming layers, resulting in $64\times 64\times 8$ subdomains. Each subdomain is then converted into a graph, where voxels become nodes connected to their adjacent neighbors. We scale both the training and the inference methods with multi-GPU pipelines. Using four RTX4090 GPUs, training takes 52hr, while inference takes 480s. Figure 5 shows the GABI-ABC pressure and velocity results.
@@ Appendix A Proofs
We remember
$Z(y)=\int_{\mathcal{U}}\exp\left(-\Phi(u;y)\right)\mathrm{d}\mathbb{P}_{u},\quad\quad\mathbb{P}_{u}:=g_{\,\#}\mathbb{P}_{z}.$
@@
where the second equality is given by Theorem 3.6.1 in Bogachev & Ruas (2007). Applying the same equality, $\forall A\in\mathcal{F}$,
@@
This follows directly from Lemma 2.1 and Theorem 2.2 by identifying $z$ with $(\mathbf{z},\sigma_{o})$, $\mathbb{P}_{z}$ with $(\mathbb{P}_{\mathbf{z}}\otimes\mathbb{P}_{\sigma_{o}})$, $g$ with $(D^{\psi}_{o}\otimes\mathrm{Id}.)$, changing to potential function to $\Phi(\mathbf{u}_{o},\sigma_{o};\mathbf{y}_{o})=\frac{1}{2\sigma_{o}^{2}}\|\mathbf{y}_{o}-\mathbf{H}_{o}\bf{\mathbf{u}}_{o}\|^{2}_{2}$, and finally noting $\mathrm{Id}._{\#}\mathbb{P}_{\sigma_{o}}=\mathbb{P}_{\sigma_{o}}$. ∎
@@ Appendix B Additional Numerical Results and Details
  • Geometry aware architecture and data-structure. As the main premise of this work is to share information about physical fields which depend on the geometry of the problem to which they relate, it is essential that the data-structure reflect this geometric information, that the neural network architecture be effective at processing it, and that both naturally accommodate inputs of varying size and topology.
  • Mapping to and from fixed-dimensional vectors. This framework relies on learning latent representations between solution to PDEs and the geometries on which these are defined. Hence, one needs to map variable-sized geometric inputs to and from fixed dimensional latent spaces.
@@ B.1 Graph Convolutional Network
In GABI: In the encoder $E^{\theta}$ we map attributed graphs to latent vectors of fixed dimensions. The decoder maps a pairing between a latent variable and a graph and populates the node attributes on the input graph. Hence the last layer in a channel wise averaging across the graph and the channel dimension is taken to be the latent dimension. The decoder $D^{\psi}$ takes as argument a graph where the node attributes are only the coordinates, hence only contains geometric information, and we expand the latent sample $\mathbf{z}$ to decode to populate all nodes with a channel dimension equal to the latent dimension.
@@ B.3.2 Additional Results
In Figure 6 we show the results of inference with GABI-ABC on for the full field reconstructing from the small number of sparse noisy observations in red. In Figure 7 we show the two terms in the loss for training the GABI autoencoder as well as a histogram of the encoded data batch overlaid with the 1D standard normal distribution. We then test the performance of ABC sampling vs NUTS sampling as we vary the dimension of the latent space for the heat inversion task. The MAE, quality of UQ, and inference times are shown in Figure 8. In Figure 9 we perform an ablation study on the size of the channel space of the graph autoencoder for a fixed latent dimension of 100 for Bayesian inversion with both ABC and MCMC sampling. From this, we can assess the importance of network size for accuracy and efficiency.
@@
Figure 8: Sampling schemes on the same autoencoder model with the same channel space and different latent dimensions. We compare (a) mean absolute predictive error, (b) uncertainty quantification under a Gaussian assumption, (c) sampling time.
(a) MAE
(b) UQ
(c) Inference time
Figure 9: Sampling schemes on the same autoencoder model with different channel space dimensions and same the latent dimension (fixed at 100). We compare (a) mean absolute predictive error, (b) uncertainty quantification under a Gaussian assumption, (c) sampling time.
@@ B.4.2 Additional Results
Figure 12: Airfoil field samples drawn from the geometry conditioned joint prior over pressure, $v_{x}$, and $v_{y}$ generated as $\mathbf{u}_{i}\sim D^{\psi}_{n\,\#}q_{\mathbf{z}}$
We now test the behaviour of GABI-ABC inference on 10 out-of-distribution airfoil geometries. In Figure 13 we plot a histogram of the density of the training dataset geometries and a selection of out of distribution airfoils. The x-axis shows the Mahalanobis distance (empirical covariance weighted distance of a point to the mean) between these shapes and the training dataset. The distance in geometry is computed using 9 summary statistics regarding airfoil thickness, camber, trailing-edge thickness, etc. We report the MAE and percentage of the ground truth between one and two standard deviations of the predictive posterior for pressure (reconstructed with GABI using 20 observations). These tests suggest that the GABI framework has some mild, but limited, robustness to prediction for outliers. Eventually, the miss-specification of the prior for such out-of-distribution geometries may lead to unreliable inference.
Figure 13: Out of distribution airfoil shapes and corresponding GABI inversion MAE and UQ for pressure, based on 20 observed airfoil nodes.